# Unveiling the Unexplored Decay Mode of a Light Charged Higgs Boson to an Off-Shell Top Quark and a Bottom Quark

### v1

@@ III $H^{\pm}\rightarrow t^{*}b$ at the HL-LHC and a 100 TeV $pp$ collider
In Figure 5, we present the normalized distributions of $\mathcal{A}_{M}$ (left) and the invariant mass $M(j_{1}j_{2}b_{1}b_{2})$ (right) for the signal with $M_{H^{\pm}}=150{\;{\rm GeV}}$ (blue) and the $t\bar{t}$ background (orange), after imposing the $\Delta R$ cuts in Equation 14. Two distinct features are evident. First, as expected, the $\mathcal{A}_{M}$ distribution for the signal peaks sharply at $\mathcal{A}_{M}\simeq 0$, while the background prefers larger values above 0.2. Imposing an upper bound on $\mathcal{A}_{M}$ will efficiently separate the signal events from the background.
@@
Table 1: Cut-flow for the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ with $M_{H^{\pm}}=150{\;{\rm GeV}}$ and the background process $pp\rightarrow t\bar{t}\rightarrow jjbb\tau\nu$. The significance is calculated considering a 10% background uncertainty and an integrated luminosity of $3{\;{\rm ab}^{-1}}$.
To evaluate the discovery potential for this channel, we present in Table 1 the cut-flow for the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ at the 14 TeV LHC. The table shows the cross sections of the signal and background after each cut, as well as the significance considering a 10% background uncertainty for an integrated luminosity of $3{\;{\rm ab}^{-1}}$. The significance is defined as:
$\mathcal{S}=\left[2(N_{\rm s}+N_{\rm b})\log\left(\frac{(N_{\rm s}+N_{\rm b})(N_{\rm b}+\delta_{\rm bg}^{2})}{N_{\rm b}^{2}+(N_{\rm s}+N_{\rm b})\delta_{\rm bg}^{2}}\right)-\frac{2N_{\rm b}^{2}}{\delta_{b}^{2}}\log\left(1+\frac{\delta_{\rm bg}^{2}N_{\rm s}}{N_{\rm b}(N_{\rm b}+\delta_{\rm bg}^{2})}\right)\right]^{1/2},$ (17)
@@
Despite applying the key kinematic cuts, the final significance reaches only 0.19. While our final selection cut boosts the signal significance by approximately 300-fold relative to the basic selection phase, the significance remains substantially below a detectable level. Moreover, the signal cross section after the final selection, approximately $4.65\times 10^{-3}{\;{\rm fb}}$, precludes any further refinements through cuts. This cut-based analysis highlights the challenges in probing the charged Higgs boson through the $H^{+}H^{-}\rightarrow t^{*}b\tau\nu$ channel at the HL-LHC using kinematic cuts alone, due to the extremely small signal cross section and overwhelming backgrounds.
To assess if the BDT method can enhance sensitivity to the $t^{*}b$ decay mode of the light charged Higgs boson, we employed the Extreme Gradient Boosting (XGBoost) package Chen:2016btl . XGBoost has seen increasing use in the particle physics community for a variety of analyses, including studies on the SM Higgs boson ATLAS:2017ztq ; CMS:2020tkr ; ATLAS:2021ifb ; CMS:2020cga ; CMS:2021sdq , dark matter ATLAS:2021jbf , vectorlike quarks Dasgupta:2021fzw , a composite pseudoscalar Cornell:2020usb , and innovative strategies for faster event generation Bishara:2019iwh . In our study, XGBoost was used as a binary classifier, aimed at more effectively distinguishing between signal and background events.
We initialized the XGBoost classifier with the objective set to binary and the evaluation metric set to logloss. The learning rate was configured at 0.1. For training the model, we generated $3.7\times 10^{4}$ signal events and $3.3\times 10^{5}$ background events, all of which met the basic selection criteria. We divided the dataset into three parts: 50% for training, 20% for validating the algorithm, and 30% for testing. As inputs for the XGBoost model, we used the following 50 variables:
@@
    Angular distance: $\Delta R(b_{1},b_{2})$, $\Delta R(b_{1},j_{1}j_{2})$, $\Delta R(b_{1},j_{1}j_{2}b_{2})$, $\Delta R(b_{1},{\tau_{\rm h}})$, $\Delta R(b_{1}{\tau_{\rm h}},j_{1}j_{2})$, $\Delta R(b_{2},j_{1}j_{2})$, $\Delta R(b_{2},j_{1}j_{2}b_{1})$, $\Delta R(b_{2},{\tau_{\rm h}})$, $\Delta R(b_{2}{\tau_{\rm h}},j_{1}j_{2})$, $\Delta R(j_{1},b_{1})$, $\Delta R(j_{1},b_{2})$, $\Delta R(j_{1},j_{2})$, $\Delta R(j_{1},{\tau_{\rm h}})$, $\Delta R(j_{2},b_{1})$, $\Delta R(j_{2},b_{2})$, $\Delta R(j_{2},{\tau_{\rm h}})$, $\Delta R({\tau_{\rm h}},j_{1}j_{2})$, $\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{1})$, $\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{1}b_{2})$, $\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{2})$.
@@
    Missing transverse energy and Missing energy azimuthal angle: $E_{T}^{\rm miss}$, $\phi^{\rm miss}$.
@@
Here, the multiple particle symbol represents the system consisting of the constituent particles. For example, the term $j_{1}j_{2}$ refers to a system whose momentum is the vector sum of the momenta of $j_{1}$ and $j_{2}$.
@@
In Figure 6, we illustrate the distributions of BDT scores for both the signal (in blue) and the background (in red). These results are derived exclusively from the testing dataset, which the model did not encounter during its training and validation phases. The BDT score distributions reveal a discernible separation between the signal and background. Applying a threshold of 0.98 for the XGBoost score results in cross sections of $1.42\times 10^{-2}{\;{\rm fb}}$ for the signal and $8.53\times 10^{-2}{\;{\rm fb}}$ for the background. Considering a 10% background uncertainty and an integrated luminosity of $3{\;{\rm ab}^{-1}}$, the signal significance reaches 1.35.
@@
In conclusion, high-energy hadron colliders, such as the HL-LHC and a prospective 100 TeV $pp$ collider, cannot effectively probe the $t^{*}b$ decay mode of the light charged Higgs boson due to the inherent softness of the $b$-jets in this channel. This limitation highlights the need for alternative approaches or collider technologies to investigate this particular decay mode of the light charged Higgs boson.
IV $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
In the previous section, we demonstrated that the proposed signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu$ at the HL-LHC yields very low significance, reaching only 1.35 even with BDT analysis. This low significance is primarily due to the $b$ jets from the decay $H^{\pm}\rightarrow t^{*}b$ being too soft for effective reconstruction. Additionally, a 100 TeV $pp$ collider cannot adequately boost the $b$ jets as the beam energy is not fully transferred to the parton-parton collision. Consequently, we shift our focus to a multi-TeV MuC, which fully harnesses the beam energy in collisions between two fundamental particles. We target two configurations: $\sqrt{s}=3{\;{\rm TeV}}$ with an integrated luminosity of $1{\;{\rm ab}^{-1}}$ and $\sqrt{s}=10{\;{\rm TeV}}$ with an integrated luminosity of $10{\;{\rm ab}^{-1}}$ Han:2021udl ; AlAli:2021let .
@@ IV $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
The first process in Equation 19 is the Drell-Yan process mediated by the $Z$ boson and photon. The second process in Equation 20 involves $H^{\pm}$ pair production associated with a pair of neutrinos. Since these additional neutrinos manifest as missing transverse energy, they create the same phenomenological signature as the final state with a single neutrino, making it indistinguishable from the first process. For this process, there are numerous Feynman diagrams, including those involving $Z^{*}/\gamma^{*}\rightarrow H^{+}H^{-}$ and $Z^{*}\rightarrow\nu\bar{\nu}$. Additionally, Vector Boson Scattering (VBS) processes such as $W^{+}W^{-}\rightarrow\gamma^{*}/Z^{*}/h^{*}\rightarrow H^{+}H^{-}$, VBS processes through the quartic vertex $W^{+}$-$W^{-}$-$H^{+}$-$H^{-}$ in Equation 7, and VBS processes through the $t$-channels mediated by $H$ and $A$ also contribute.
@@
The third process in Equation 21 involves $H^{\pm}$ pair production associated with two forward muons. In the conventional detector design of a multi-TeV MuC, forward muons are not detectable [^3] Recently, the integration of forward muon detectors into the MuC’s design has been initiated within the MuC community, as the energetic forward muons can penetrate the tungsten nozzles. This forward muon detector is highly expected to play a crucial role in probing various BSM models Accettura:2023ked ; Ruhdorfer:2023uea ; Forslund:2023reu ; Bandyopadhyay:2024plc . because they fall outside the pseudorapidity coverage limit of $|\eta|<2.5$. This limitation is due to tungsten nozzles designed to shield the detector from BIB particles Collamati:2021sbv ; Ally:2022rgk . This neutral-current VBS process occurs through $ZZ\to h^{*}\to H^{+}H^{-}$, the quartic vertex $Z/\gamma$-$Z/\gamma$-$H^{+}$-$H^{-}$, and the $t$-channels mediated by $H$ and $A$.
@@
Figure 7: Parton-level cross sections as a function of $M_{H/A}$ for the processes ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ (left) and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ (right) for $M_{H^{\pm}}=150{\;{\rm GeV}}$ at c.m. energy $\sqrt{s}=10{\;{\rm TeV}}$, over the allowed parameter points. $\mu_{\rm f}$ denotes the forward muons with $|\eta_{\mu}|>2.5$. The color scale represents the values of $\tan\beta$.
@@
Let us discuss the dependence of cross sections for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ on model parameters. Unlike the Drell-Yan process, these two associated production processes include contributions mediated by $H$ and $A$, making their cross sections potentially sensitive to $M_{H/A}$. Additionally, VBS contributions through $W^{+}W^{-}\rightarrow h^{*}\rightarrow H^{+}H^{-}$ to ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and $ZZ\to h^{*}\to H^{+}H^{-}$ to ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$, though not dominant, introduce dependences on $\tan\beta$.
@@
Figure 7 illustrates parton-level cross sections as a function of $M_{H/A}$ for the processes ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ with $M_{H^{\pm}}=150{\;{\rm GeV}}$ at $\sqrt{s}=10{\;{\rm TeV}}$. The color codes denote $\tan\beta$. The analysis considers parameter points allowed by theoretical and experimental constraints discussed in Sec. II. For forward muons $\mu_{\rm f}$, we apply the condition $4.74132.5<|\eta_{\mu}|<$ 4.7413. The lower bound of $|\eta|>2.5$ ensures that forward muons fall outside the conventional pseudorapidity coverage limit. The upper bound of $|\eta|<4.7413$ is implemented to manage cross section divergence that occurs as the scattering angle $\theta$ approaches 0 or $\pi$, a consequence of $t$-channel photon-mediated diagrams. Typically, this divergence is canceled by higher-order QED corrections, particularly from soft photon emissions. In practice, such divergences are often handled by imposing a minimum scattering angle cut during data analysis. Considering the high collision energy at the MuC, we constrain the scattering angle to $1^{\circ}<\theta<179^{\circ}$, which corresponds to $|\eta|<4.7413$.
@@
Figure 7 clearly demonstrates that both processes exhibit modest dependence of cross sections on the model parameters $M_{H/A}$ and $\tan\beta$. Notably, the cross section for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ remains nearly constant across the allowed parameter space, varying by only about 10%. In contrast, the cross section for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ shows a more pronounced dependence on the model parameters, with variations of approximately 70%. For a given $\tan\beta$, the cross section initially decreases as $M_{H/A}$ increases, reaches a minimum at $M_{H/A}\simeq 200{\;{\rm GeV}}$, and then rises again.
@@
Given these observations, we adopt a conservative approach by selecting $M_{H}=M_{A}=200{\;{\rm GeV}}$, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}{\;{\rm GeV}}^{2}$ for our signal-to-background analysis. This choice represents a pessimistic scenario, ensuring that if we achieve signal significance above the discovery potential, it would guarantee the accessibility of the entire parameter space through our target signal $H^{\pm}\to t^{*}b$.
@@
We now compare the cross sections of the three production channels for the charged Higgs mass:

### v2

@@ 3 $H^{\pm}\rightarrow t^{*}b$ at the HL-LHC and a 100 TeV $pp$ collider
In Figure 5, we present the normalized distributions of the asymmetry variable $\mathcal{A}_{M}$ (left panel) and the invariant mass $M(j_{1}j_{2}b_{1}b_{2})$ (right panel) for the signal with $M_{H^{\pm}}=150{\;{\rm GeV}}$ (blue lines) and the $t\bar{t}$ background (orange histograms), after imposing the $\Delta R$ cuts defined in Equation 33. Two distinct features are immediately evident. First, as expected, the $\mathcal{A}_{M}$ distribution for the signal peaks sharply near $\mathcal{A}_{M}\simeq 0$, whereas the background distribution is concentrated at higher values, typically above 0.2. Thus, applying an upper bound on $\mathcal{A}_{M}$ will effectively enhance signal-background separation.
@@
Table 1: Cut-flow for the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ with $M_{H^{\pm}}=150{\;{\rm GeV}}$ and the dominant background process $pp\rightarrow t\bar{t}\rightarrow jjbb\tau\nu$. The significance is calculated considering a 10% background uncertainty and an integrated luminosity of $3{\;{\rm ab}^{-1}}$.
To evaluate the discovery potential for this channel, we present in Table 1 the cut-flow for the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow jjbb\tau\nu$ at the 14 TeV LHC. The table shows the cross sections of the signal and background after each cut, as well as the significance considering a 10% background uncertainty for an integrated luminosity of $3{\;{\rm ab}^{-1}}$. The significance is defined as Cowan:2010js :
$\mathcal{S}$ $=$ $\left[2(N_{\rm s}+N_{\rm b})\log\left(\frac{(N_{\rm s}+N_{\rm b})(N_{\rm b}+\delta_{\rm bg}^{2})}{N_{\rm b}^{2}+(N_{\rm s}+N_{\rm b})\delta_{\rm bg}^{2}}\right)\right.$
$\left.-\frac{2N_{\rm b}^{2}}{\delta_{b}^{2}}\log\left(1+\frac{\delta_{\rm bg}^{2}N_{\rm s}}{N_{\rm b}(N_{\rm b}+\delta_{\rm bg}^{2})}\right)\right]^{1/2},$
@@
Despite applying the key kinematic cuts, the final significance reaches only 0.19. While our final selection cut boosts the signal significance by approximately 300-fold relative to the basic selection step, the significance remains substantially below a detectable level. Moreover, the signal cross section after the final selection, approximately $4.65{\;{\rm ab}}$, precludes any further refinements through cuts. This cut-based analysis highlights the challenges in probing the charged Higgs boson through the $H^{+}H^{-}\rightarrow t^{*}b\tau\nu$ channel at the HL-LHC using kinematic cuts alone, due to the extremely small signal cross section and overwhelming backgrounds.
To assess if the BDT method can enhance sensitivity to the $t^{*}b$ decay mode of the light charged Higgs boson, we employed the Extreme Gradient Boosting (XGBoost) package Chen:2016btl . XGBoost has seen increasing use in the particle physics community for a variety of analyses, including studies on the SM Higgs boson ATLAS:2017ztq ; CMS:2020tkr ; ATLAS:2021ifb ; CMS:2020cga ; CMS:2021sdq , dark matter ATLAS:2021jbf , vectorlike quarks Dasgupta:2021fzw , a composite pseudoscalar Cornell:2020usb , and innovative strategies for faster event generation Bishara:2019iwh . In our study, XGBoost is used as a binary classifier, aimed at more effectively distinguishing between signal and background events.
We initialized the XGBoost classifier with the objective set to binary and the evaluation metric set to logloss. The learning rate was configured at 0.1. For training the model, we generated $3.7\times 10^{4}$ signal events and $3.3\times 10^{5}$ background events, all of which met the basic selection criteria. We divided the dataset into three parts: 50% for training, 20% for validating the algorithm, and 30% for testing. As inputs for the XGBoost model, we used the following 51 variables: [^5] A large number of input parameters does not pose a concern in XGBoost, as its tree-based structure inherently selects only the most relevant features. This prevents overfitting and ensures that unnecessary or correlated inputs do not impact model performance.
@@
    Angular distance:
    $\Delta R(b_{1},b_{2}),\quad\Delta R(b_{1},j_{1}j_{2}),\quad\Delta R(b_{1},j_{1}j_{2}b_{2}),$
    $\Delta R(b_{1},{\tau_{\rm h}}),\quad\Delta R(b_{1}{\tau_{\rm h}},j_{1}j_{2}),\quad\Delta R(b_{2},j_{1}j_{2}),$
    $\Delta R(b_{2},j_{1}j_{2}b_{1}),\quad\Delta R(b_{2},{\tau_{\rm h}}),\quad\Delta R(b_{2}{\tau_{\rm h}},j_{1}j_{2}),$
    $\Delta R(j_{1},b_{1}),\quad\Delta R(j_{1},b_{2}),\quad\Delta R(j_{1},j_{2}),$
    $\Delta R(j_{1},{\tau_{\rm h}}),\quad\Delta R(j_{2},b_{1}),\quad\Delta R(j_{2},b_{2}),$
    $\Delta R(j_{2},{\tau_{\rm h}}),\quad\Delta R({\tau_{\rm h}},j_{1}j_{2}),\quad\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{1}),$
    $\Delta R({\tau_{\rm h}},j_{1}j_{2}),\quad\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{1}b_{2}),\quad\Delta R({\tau_{\rm h}},j_{1}j_{2}b_{2}).$
@@
    Missing transverse energy and Missing energy azimuthal angle:
    $E_{T}^{\rm miss}$, $\phi^{\rm miss}$.
@@
Here, a symbol representing multiple particles denotes a composite system formed by its constituent particles. For example, in $\Delta R({\tau_{\rm h}},j_{1}j_{2})$, the term $j_{1}j_{2}$ represents a system composed of $j_{1}$ and $j_{2}$, where the four-momentum of this composite system is defined as the vector sum of the individual momenta of $j_{1}$ and $j_{2}$. Thus, $\Delta R({\tau_{\rm h}},j_{1}j_{2})$ corresponds to the angular distance between ${\tau_{\rm h}}$ and the combined system of $j_{1}$ and $j_{2}$. This convention is consistently applied throughout our feature set.
@@
In Figure 6, we illustrate the distributions of BDT scores for both the signal (blue) and the background (red). These results are derived exclusively from the testing dataset, which the model did not encounter during its training and validation phases. The BDT score distributions reveal a discernible separation between the signal and background. Applying a threshold of 0.98 for the XGBoost score results in cross sections of $1.42\times 10^{-2}{\;{\rm fb}}$ for the signal and $8.53\times 10^{-2}{\;{\rm fb}}$ for the background. Considering a 10% background uncertainty and an integrated luminosity of $3{\;{\rm ab}^{-1}}$, the signal significance reaches 1.35.
@@
In conclusion, high-energy hadron colliders, such as the HL-LHC and a prospective 100 TeV $pp$ collider, cannot effectively probe the $t^{*}b$ decay mode of the light charged Higgs boson due to the inherent softness of the $b$-jets in this channel. This limitation highlights the need for an alternative collider to investigate this particular decay mode of the light charged Higgs boson.
4 $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
Our analysis in the previous section showed that the signal process $pp\rightarrow H^{+}H^{-}\rightarrow t^{*}b\tau\nu$ yields insufficient significance at hadron colliders, primarily because the $b$ jets from $H^{\pm}\rightarrow t^{*}b$ decay are too soft for effective reconstruction. Even at a 100 TeV $pp$ collider, where only a fraction of the beam energy contributes to the parton-parton collision, the $b$ jets remain insufficiently boosted. This limitation motivates us to consider a multi-TeV MuC, where the full beam energy is available in collisions between fundamental particles. We examine two configurations Han:2021udl ; AlAli:2021let : $\sqrt{s}=3{\;{\rm TeV}}$ with an integrated luminosity of $1{\;{\rm ab}^{-1}}$ and $\sqrt{s}=10{\;{\rm TeV}}$ with $10{\;{\rm ab}^{-1}}$.
@@ 4 $H^{\pm}\rightarrow t^{*}b$ at a Multi-TeV Muon Collider
The first process in Equation 38 represents the Drell-Yan process mediated by the $Z$ boson and photon. The second process in Equation 39 involves $H^{\pm}$ pair production with two additional neutrinos. Since these neutrinos appear as missing transverse energy in the detector, this process creates a signature identical to our target process with a single neutrino in Equation 37. For this process, there are multiple contributing Feynman diagrams, with vector boson scattering (VBS) processes through $t$-channel $H$ and $A$ mediators being particularly important. We provide a detailed discussion of this process in Appendix A.
@@
The third process in Equation 40 describes $H^{\pm}$ pair production with two forward muons. In conventional multi-TeV MuC detector designs, these forward muons cannot be detected due to tungsten nozzles [^6] Recently, the MuC community has begun integrating forward muon detectors into their designs, as energetic forward muons can penetrate the tungsten nozzles. These detectors are expected to play a crucial role in probing various BSM models Accettura:2023ked ; Ruhdorfer:2023uea ; Forslund:2023reu ; Bandyopadhyay:2024plc . that shield the detector from BIB particles Collamati:2021sbv ; Ally:2022rgk , limiting pseudorapidity coverage to $|\eta|<2.5$. This neutral-current VBS process proceeds through several channels: $ZZ\to h^{*}\to H^{+}H^{-}$, the quartic vertex $Z/\gamma$-$Z/\gamma$-$H^{+}$-$H^{-}$, and $t$-channel exchanges of $H$ and $A$.
@@
Figure 7: Parton-level cross sections as a function of $M_{H/A}$ for the processes ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ for $M_{H^{\pm}}=130{\;{\rm GeV}}$ (upper) and $M_{H^{\pm}}=150{\;{\rm GeV}}$ (lower) at c.m. energy $\sqrt{s}=10{\;{\rm TeV}}$, over the allowed parameter points. $\mu_{\rm f}$ denotes the forward muons with $|\eta_{\mu}|>2.5$. The color scale represents the values of $\tan\beta$.
@@
Let us discuss the dependence of cross sections for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ on model parameters. Unlike the Drell-Yan process, these two processes include contributions mediated by $H$ and $A$, making their cross sections potentially sensitive to $M_{H/A}$. Additionally, VBS contributions through $W^{+}W^{-}\rightarrow h^{*}\rightarrow H^{+}H^{-}$ to ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and $ZZ\to h^{*}\to H^{+}H^{-}$ to ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ introduce dependences on $\tan\beta$.
Figure 7 illustrates the parton-level cross sections as a function of $M_{H/A}$ for the processes ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ with $M_{H^{\pm}}=130~{}{\;{\rm GeV}}$ (upper panel) and $M_{H^{\pm}}=150~{}{\;{\rm GeV}}$ (lower panel) at $\sqrt{s}=10~{}{\;{\rm TeV}}$. The color codes denote $\tan\beta$. The analysis considers parameter points allowed by theoretical and experimental constraints discussed in Sec. 2. The cross sections were computed using MadGraph5_aMC@NLO Alwall:2011uj , which includes all relevant tree-level Feynman diagrams contributing to these $2\to 4$ processes.
@@
For the process involving forward muons ($\mu_{\rm f}$), we apply the condition $4.74132.5<|\eta_{\mu}|<$ 4.7413. The lower bound ensures that forward muons fall outside the conventional pseudorapidity coverage limit of the MuC. The upper bound, $|\eta|<4.7413$, is introduced to regulate the divergence that occurs as the scattering angle $\theta$ approaches 0 or $\pi$, a consequence of $t$-channel photon-mediated diagrams. This divergence is typically canceled by higher-order QED corrections, particularly from soft photon emissions. In practice, such effects are handled by imposing a minimum scattering angle cut during data analysis. Considering the high collision energy at the MuC, we apply the angular constraint $1^{\circ}<\theta<179^{\circ}$, which corresponds to $|\eta|<4.7413$.
Figure 7 shows that the results for $M_{H^{\pm}}=130{\;{\rm GeV}}$ exhibit similar trends to those for $M_{H^{\pm}}=150{\;{\rm GeV}}$, with two notable differences. First, the cross sections for $M_{H^{\pm}}=130{\;{\rm GeV}}$ are moderately larger than those for $M_{H^{\pm}}=150{\;{\rm GeV}}$, increasing by approximately 20%. This enhancement is due to the lighter $M_{H^{\pm}}$. Second, the upper bounds on $M_{H/A}$ are lower for $M_{H^{\pm}}=130{\;{\rm GeV}}$ than for $M_{H^{\pm}}=150{\;{\rm GeV}}$, as dictated by the theoretical constraints of perturbativity and unitarity, which prohibit excessively large mass differences between $M_{H^{\pm}}$ and $M_{H/A}$ (see Equation 23).
@@
Another important observation from Figure 7 is that the cross sections for both ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ exhibit only a mild dependence on the model parameters $M_{H/A}$ and $\tan\beta$. Notably, the cross section for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ remains nearly constant across the allowed parameter space, varying by only about 10%. In contrast, the cross section for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ shows a more pronounced dependence on the model parameters, with variations of approximately 50%. For a given $\tan\beta$, the cross section initially decreases as $M_{H/A}$ increases, reaches a minimum at $M_{H/A}\simeq 182~{}(200){\;{\rm GeV}}$ for $M_{H^{\pm}}=130~{}(150){\;{\rm GeV}}$, and then rises again. A detailed discussion of the dominant Feynman diagrams and the dependence of $\sigma({\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu})$ on model parameters can be found in Appendix A.
@@
Given these observations, we adopt a conservative approach by selecting $M_{H}=M_{A}=200{\;{\rm GeV}}$, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}{\;{\rm GeV}}^{2}$ for our signal-to-background analysis. This parameter point represents a pessimistic scenario, yielding the lowest cross section for ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ within the allowed parameter space with $M_{H^{\pm}}=150{\;{\rm GeV}}$. If we can achieve a discovery-level signal significance for this conservative benchmark, parameter points with higher cross sections would provide better detection prospects.
Nevertheless, we note that the discovery reach could be moderately affected by different choices of model parameters through small variations in kinematics and event selection efficiencies. Such effects are more relevant at the 10 TeV MuC, where the cross section of ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ is comparable to that of ${\mu^{+}\mu^{-}}\to H^{+}H^{-}$. In contrast, at the 3 TeV MuC, the cross section of ${\mu^{+}\mu^{-}}\to H^{+}H^{-}$ is an order of magnitude larger than those of ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$ and ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\mu_{\rm f}^{+}\mu_{\rm f}^{-}$ (these cross sections will be presented in Table 2). As a result, the discovery potential at 3 TeV exhibits much weaker dependence on the model parameters, since the ${\mu^{+}\mu^{-}}\to H^{+}H^{-}$ process is independent of these parameters for a given $M_{H^{\pm}}$.
@@
We now compare the cross sections of the three production channels in Equation 38, Equation 39, and Equation 40 for the charged Higgs mass